State and derive the ohms law?                                      😁😁😁😁😁😁😁😁😁😁😁

Grade:10

2 Answers

Arun
4 years ago
Ohm' s law :- the current flowing through a conductor is directly proportional to the potential difference across it ends.

where ,
V = pd.
R = resistanceΒ 
I =elec. currentΒ 

therefore ,
V = RIΒ 
R = V / I

V / I = constant
therefore ,
I = resistance
Β 
Β 
Khimraj
4 years ago

Statement:– Ohm’s law states that the current (I) flowing through a conductor is directly proportional to the potential difference (V) across the ends of the conductor,provided physical conditions of the conductors such as temperature,mechanical strain etc. Are kept constant i.e. IΞ±V

OrΒ Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β  VΞ±I or V= RI

OrΒ Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β  V/I=R=a constant

Where R is known as resistance of the conductor . It depends upon the length ,shape and the nature of the material of the conductor. The variation between potential difference (V) and current (I) through a conductor is a straight line.

Deduction of Ohm’s Law:

AsΒ drift velocity isΒ given by Vd=eEt/m

ButΒ Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β  Electric field across the conductor of length l is

E=V/l

ThusΒ Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β  Vd=e V/ml

Also,Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β  I=Anevd

By putting value of vdΒ in the equation of I ,we get

I=Ane(eV/ml)=(Ane2/ml)V

OrΒ Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β  V/I=ml/Ane2=R

Where R is a constant for a given conductor. It is known as the resistance of the conductor.

ThusΒ Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β Β  V=RI

ThisΒ  is Ohm’s law.